OpenStudy (anonymous):

x and y vary inversely, x=50 when y=5. what is value of k?

OpenStudy (anonymous):

x=Ky so we know that 50=K*5 inverse operations from there

OpenStudy (anonymous):

sorry I just noticed inversely

OpenStudy (anonymous):

x=K/y 50=K/5 inverse operations from there

OpenStudy (anonymous):

i came up with 250?

OpenStudy (anonymous):

sounds legit